OpenStudy (anonymous):

Find a number k such that the given equation has exactly one real solution. x^2-kx+4=0

myininaya (myininaya):

k=4 since x^2-4k+4=(x-2)^2=0 which gives solution x=2 multiplicity 2

myininaya (myininaya):

or you might like this better...

myininaya (myininaya):

to complete the square we do \[(-k/2)^2\] but we want this to be 4 so set them equal and solve for k

myininaya (myininaya):

\[(\frac{-k}{2})^2=4\]

myininaya (myininaya):

\[\frac{-k}{2}= \pm 2 => -k =\pm 4 => k= \mp 4\]

myininaya (myininaya):

so -4 would also have worked